OpenStudy (anonymous):

What would a sub 7 equal, if a sub 1 equals 8 and a sub k+1 equals a sub k plus 3?

OpenStudy (anonymous):

Let \[ a _{1}=8\] and \[a _{k+1}=a _{k}+3\] Determine \[a_{7}\]

OpenStudy (akashdeepdeb):

This is known as a recursive function. We are given that \(a_1 = 8\) So How much does \(a_2\) equal? Well we know that, \(a_{k+1} = a_k + 3\) So, \(a_2 = a_{1+1} = a_1 + 3 = 8 + 3 = 11\) [Here k becomes 1] Hence, now we have the value of \(a_2 = 11\). Did you get this recursion? Can you proceed? :)

OpenStudy (jdoe0001):

\(\large { a_1=8\qquad a_{k+1}=a_k+3 \\ \quad \\ \begin{array}{ccrcllll} term(k)&&value \\\hline\\ {\color{brown}{ 1}}&a_1&8\\ {\color{brown}{ 2}}&a_{{\color{brown}{ 1}}+1}&a_{\color{brown}{ 1}}+3\to 8+3\to 11\\ {\color{brown}{ 3}}&a_{{\color{brown}{ 2}}+1}&a_{\color{brown}{ 2}}+3\to 11+3\to 14\\ 4&a_{{\color{brown}{ 3}}+1}& a_{{\color{brown}{ 3}}}+3\to 14+3\to 17\\ ...&...&... \end{array} }\)

OpenStudy (jdoe0001):

keep in mind that the notation \(\bf a_{k+1}\) just means, " the next term " so \(\bf a_{k+1}=a_k+3\to \) whatever the "current term" is, +1, or next term, will be our "current value" + 3

OpenStudy (anonymous):

oh okay! thank you both so much! @AkashdeepDeb @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (akashdeepdeb):

:)